12. What is the hydroxide ion concentration of bile (pH = 8.4)?! POH Answer: [OH-] = 2.5IXOS..
Answer #1

As we know, pH tpot = 14. In Given problem pH = 8.4 This poH = 14-8.4. - 5.6 By deffination port = -log[64] » [oH] = 10-56 »